OpenStudy (anonymous):

what does this mean? 9x - 7i > 3(3x -7u) will give medal and fan!

OpenStudy (anonymous):

Well i don't know what that means

OpenStudy (anonymous):

haha

OpenStudy (anonymous):

Instead of writing y = 3x + 2 you may write f(x) = 3x + 2 or g(x) = 3x + 2 or perhaps even y(x) = 3x + 2

OpenStudy (anonymous):

it still don't make since

OpenStudy (destinymasha):

;)

OpenStudy (anonymous):

will some one tell us?

OpenStudy (anonymous):

i think no one will what u wanna know...everyone sees this as a maths que

OpenStudy (anonymous):

crap thats right